Sol Summary – December 4th

MDRS Daily Summary Report for Sol 3

Summary Title: First full up day of activities!

Mission Status: Today really felt like our first day on mars. It was our first full activity day as well as our first day on EVA! Surprisingly it was consensus among our crew that our first full day was actually the easiest day we’ve had so far! I think the excitement of going outside and our team getting into the swing of things has really made us more efficient.

The primary activity for the day was an EVA on all crew to familiarize ourselves with EVA procedures, suit functionality, and operation of the rovers. Many lessons were learned, and it was good that we got the kinks out on a EVA solely suited to it. This will allow our scientific EVAs to be successful later on, as well as ensure that we can safely perform EVAs at a farther distance from the habitat.

Our team is following a specific protocol for depress / repress operations modeled after ISS EVAs. Our team is doing a 25 minute prebreathe in the EVA Prep room to prevent decompression sickness, once complete we are doing a 5 minute depressurization in the actual airlock. During repress we are allocating 5 minutes for repressurization options, though we have the capability to do an ’emergency repress’ if needed at the risk of rupturing crew eardrums.

Below are the several lessons learned we had during EVA.

  1. Peripheral vision is nearly non-existent when wearing the EVA helmets. This makes it difficult to travel together in an ATV convoy as it is impossible to turn around and check on the status of those behind you. We attempted to mitigate this by placing the radios in VOX and keeping in constant voice contact to maintain proper spacing, we also played ‘leapfrog’ with the ATV’s where the leader would stop after several minutes of driving to allow the two behind him to catch up and pass, ensuring no-one was left behind.
  2. We had several communication failures during EVA due to earpieces falling out of ears. We addressed this via using medical tape to secure the earpieces in place.
  3. We had to abort a prebreathe operation due to forgetting to bring supplies into the EVA Prep room. We instituted a supply check prior to closing the Prep Room door. If necessary to break EVA prebreathe protocol we do a 3 to 1 buy back, i.e. if you break prebreathe for 5 minutes, you add 15 minutes to prebreathe to compensate.
  4. Several of our crew members are trained scuba divers. We found that it was difficult to hear other EVA crew while driving the ATV due to engine and wind noise, we instituted a series of hand signals to allow non-verbal communication which quickly and concisely allowed EVA crew status to be conveyed.
  5. Do not wear high pony tails – recommend low braid. A high braid interferes with the way that your head can move within the EVA helmet and vision is confined to roughly a 45 degree angle down and lower.

During the high winds today we heard a bang as the outer airlock door was blown open. The inner airlock door held closed and thus we didn’t all die! We immediately closed the door to the EVA Prep room to form a second seal to the martian atmosphere. During tomorrows familiarization EVA we will enter the EVA Prep room, close the EVA Prep room door, and use that as a depressurization chamber. Immediately upon departure on tomorrows EVA we will repair the outer airlock door.

We also completed our first scientific objectives today, covered below.

Reduced Gravity CPR Experiment: When on Mars there is a possible complication when performing CPR. With your body weight reduced the typical CPR techniques are non functional due to reduced force on the recipient. Our team is practicing techniques under the supervision of our medical officer to allow force application in excess of body weight and measuring output using a springed scaled.

Communication Delay ‘Bricks’ Experiment: Our first two experiment runs of the brick experiment were performed. We started with a 0 time delay baseline, and increased to a 9 minute time delay with verbal information transfer (martian time delay). Our runs will continue to collect more data and determine the most efficient methods of information transfer.

We also completed some major maintenance tasks today

The Musk Observatory has been set up and put through it’s paces and is ready for operation.

Trailer Tank to Static Water Tank transfer operation – Crew completed the transfer this morning completely filling our static water tank. Heartfelt thanks to the Team supporting us for dropping the trailer tank and instructions on how to transfer the water. The MDRS Hab Manual was a roaring success in allowing us to complete this task!

Sol Activity Summary:

  1. Tasks mentioned above
  2. Exercise
  3. Dirt dehydration
  4. Science Dome organization and clean up
  5. Repair of Roof Hatch hole
  6. Our first shower! Heartily enjoying the new heater
  7. Completion of reports (soon)
  1. EVA Request
  2. Sol Summary
  3. Greenhab
  4. Journalist
  5. Operations
  6. Astronomer
  7. Inventory

WiFi range survey Radio check with all radios (all functional except for 1) Look Ahead Plan:

1. Matryoshka EVA #1

2. Communication ‘Lego Bricks’ experiment

3. Sleep Study / crew medical questionnaires

4. Musk Observatory Observation

5. Journalist interviews and filming

6. Medical status surveys with Crew Doctor throughout the day.

7. CPR Experiment

8. HAB Cleaning

9. Radio Range EVA – This activity is cancelled due to 1 EVA per day limit. It will be rescheduled for a future day. We will work with schedulers to ensure they understand this limit for the future.

Anomalies in work:

  1. ATV wobbly wheel
  2. Non Functional Power Strip (removed and replaced)
  3. 3 Non Functional Radios (does not impact operations)
  4. HAB paneling peeling off due to wind damag

Weather: Very mild weather. This was welcome after yesterday.

Crew Physical Status: All crew in good health.

EVA: 2 Familarization EVAs today

Support Requested:

  1. ATV Fuel
  2. Food Drop (not urgent)


Astronomy Report – December 4th

Facility Utilized: Musk Observatory

Crewmembers Utilized: Thomas Horn

Supplies Utilized: Solar Telescope / Eyepiece / Camera / Dome

Duration (minutes): 180 minutes

Description of Activity: Radiation is a deadly threat to astronauts both in transit to and on the surface of mars, particularly when the astronauts are exposed to the sun. While astronauts can rely on ground support to identify any solar hazards that develop, communication with the ground cannot be guaranteed. If a Mars Habitat had an independent method of solar observation it could provide critical advance notice that solar flares or other hazards were developing and give direction to crew to shelter in place in radiation storm shelters.

Today’s activity was primarily focused on setting up the solar observatory for operation throughout the mission. A quick survey of the sun was done and no unusual solar activity was spotted. Images below were taken during today’s operation.

Operations Report – December 4th


Name of person filing report: HuntJ

Non-nominal systems:

  1. Wi-Fi Modem (lack of range).  The wifi in a roughly 5 feet circle from the router, any farther and signal drops off line.  We have tested this and created a signal service map by carrying laptops around the habitat to determine when it picks up signal.  No known fix, will operate internet within this range for now.
  2. HAB Eastern Wall has sustained wind damage, panels are beginning to peel off wall. When wind blows the panels swing and hit support beam causing HAB to vibrate.  In order to fix we need HAB material adhesive delivered, after which we will go EVA and patch the damage,
  3. ATV Yamaha 250 #3 Back Right Wheel is loose, observed to be wobbling during driving.  We will assess a fix during tomorrows EVA – it is possible it simply needs to be tightened, if there is actual mechanical damage then we will move the ATV and declare it off limits.
  4. 3 broken old radios.  All newly delivered radios are functional (thank you supply drop).  Assessing broken radios to determine a fix, if none identified will mark as trash and place in a ‘broken equipment bin’.  4 EVA earpieces were declared nonfunctional, though this may have been due to incorrect settings, will continue to investigate during tomorrows EVA

Generator (hours run): 15.4 Hours

Solar— SOC % : @ 7:34 AM: 65%

                            @ 8:00 AM: 100 % (Generator stopped at this time)

@ 5:34 AM: 42% (Generator started at this time)

Diesel –  48% Full.

Propane –  79 Percent Volume .

Ethanol Free Gasoline (5 Gallon containers for ATV) –   gallons. 2.5 Gallons. One 5 gallon tank empty, second tank half full. Request for extra fuel.

Water (trailer) –  600 Gallons

Water (static) –   500 gallons (We have water again!!)

Trailer to Static Pump used –  Yes

Water (loft) – Static to Loft Pump used –  Yes

Water Meter:  35 gallons

Toilet tank emptied: No

ATV’s Used: 3 Yamaha 250’s and 1 Honda ATV used.

Oil Added? No

ATV Fuel Used:  7.5 gallons used to fill all ATVs to full, though the ATVs did not start with full tanks for our mission.

# Hours the ATVs were Used today: 02:30 hours

Notes on ATVs: See above for wobbly ATV wheel issue.

Deimos rover used: Not used


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used:  ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:


Spirit rover used:  Not used


Beginning charge:

Ending charge:

Currently charging:


Opportunity rover used:  Not Used


Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  Not Used


Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: Trailer to Static water tank was used today. Pump was dirty and needed cleaning and we did so. Actual process was nominal. We are now full on water again. When inspecting trailer tank water the front tank had a layer of FOD and detritus at the water line. We will keep watch of filter usage in case this increases our burn through rate on filters.

Summary of internet:  Wifi was fixed this morning, though range is limited (see above). Please send hardware manual for the router and modem to allow us to diagnose issues and repair anything that comes up in the future (for example – to allow us to diagnose what certain light combinations signify)

Summary of suits and radios: See above

Summary of Hab operations: Brick Communication Experiment / CPR experiments successful, exercise on all crew. EVA Outer Door repaired.

Summary of GreenHab operations: No fire extinguisher observed – please deliver in future supply drop.

Summary of ScienceDome operations: Soil Salinity experiments (previous crew), dehydrating dirt for use as counterbalance and weights in exercise equipment (filling bags and medicine balls)

Summary of RAMM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support:

Need extra fuel for ATVs. Of the two red fuel cans, one is empty, one is half full.

Pre-Mission food inventory is complete, will be sent as it’s own report. Please reference that report for requested supplies to address gaps in current inventory. Please give ETA for supply drop so we can adequately ration food.

Thank you for your assistance!

GreenHab Report – December 4th

GreenHab Report

Trisha Randazzo

December 4, 2017

Environmental control: (Choose which is appropriate and explain further if needed)

Ambient (no heating or/cooling)

Ambient with window/door open

  • No wind, door open starting at 11:00
  • Plants looked a bit warm so will open door a before 10:00 tomorrow


  • Functioning nominally


  • Not available

Both heating and cooling

  • Nominal

Shade cloth on/off

  • On

Average temperature: (N/A until new sensor is delivered)

  • N/A

Changes to crops: Note all emergence of seedlings, death of seedlings, etc

  • Same tomato is still doing poorly, but watering it
  • Some tomato plants were a bit dropping, possibly due to the heat. Will open Hab door earlier than 11:00

Daily water usage for crops:

  • Complete with no issues

Time(s) of watering for crops:

  • 1:00 PM

Morning research observations: (Currently not operational)

  • N/A

Changes to research plants:

  • N/A

Daily watering and amount of water used:

  • 2 gallons for all soils

Aquaponics: (Currently not operational)

  • N/A

Narrative: Any other information you want to share

  • Paper-whites are looking good!!

Support/supplies needed:

  • Awaiting:

Journalist Report – December 4th

Crew 184 Journalist Report

Willie Schumann

03 December 2017

Title First Steps

Narrative Today was the day, we walked the first time on Martian soil. But before our boots printed our profiles in the dusty ground, we had to learn how to survive. Meaning how to use our space suits properly and to stay always in touch with our crewmates.

We decided to split our crew in groups of three each, to have a short familiarization EVA. Commander Horn and Medical Officer Sczepaniak and me had the privilege to be in group one. The pole position also helped to clear some problems for the rest of the team. It took us some time to have all radios checked and synchronized and applied perfectly on our suits. The cameraman forgot to take his camera in all his euphoria into the preparation room. But there was no time for shame for yours truly, only time for EVA.

Our space helmets are certainly a piece of art. Quite unique in size and design, and clearly differentiable to other groundbreaking space missions in the past decades. Every era should have their iconic space suits and round and clear helmets define the era of humans on Mars. It takes two people two put the helmet and the oxygen container on the astronaut, which is clearly an effort. But it also increases the feeling of security; because it clearly has what the astronaut needs most – plenty of breathable air.

But it takes a little bit more to be able to go out of the habitat than just putting on a suit. To avoid the risk of decompressing our lungs we stayed twenty five minutes in the preparation room. We used this time to triple check our before we went into the airlock. This tiny chamber is the last thin border to the Martian environment. There we decompressed for another five minutes. The moment of opening this last door to adventure was beautiful, commander Horn opened the lock and the Martian sun welcomed us.

Yesterday night blew a hefty Martian storm over our habitat, so first thing we did was checking the premises for possible damage. We discovered nothing and jumped on our ATV’s, which we had extensively trained on. I had struggled with it before, but today it was a complete joy ride. Everything seemed in sync, as I was ever meant to be to ride on this surface.

Our plan was to explore the near surroundings and so we did. The ground was mainly perfectly even and we quickly distanced ourselves from our hab. The landscape got more surreal as we gained ground. Round shaped and sharp edged hills and rocks are sprinkled left and right of us. We stopped and climbed on the highest elevation point. I was surprised how easy we reached the peak. The surface is soft and gives enough grip to step up. On top it was a great opportunity to shoot some pictures to send home.

After we returned to the hab and a quick lunch I decided to join the second group for their first minutes. I thought it was a great chance to get some additional film footage. After everything was done I actually felt the exhaustion of the EVA’s. While I was out there the adrenaline kept me focused and going, because every new shot was a promise. But it was great to return to the hab, which turned to our home within just a few days. I never expected that, but the intensity of our endeavor seems to accelerate everything. Mars really makes every minute count and precious.

Personal Logbook I am exhausted, but very happy after this day with 1,5 EVA’s. While the first one was a proper one, I took the chance to film a few transition shots in the beginning of the EVA of our second group out on Mars.

There are challenges to film with the round helmet, but I surprisingly managed well to focus. It seems I am fit to film on Mars. For the first time I also had the chance to take some photography on film on Mars, which made me really content.

I see pictures and stories everywhere, but you have to stay contained and you have to shift certain scenes and topics to another day. Otherwise I would kill myself with work. I have to stay fit and focused this entire mission.

I am happy, that the crew is picking up on my idea of VLogs. Instead of an interview I would love everyone to film personal diaries with an iPhone. Other than a conducted interview a VLog gives the audience the chance to have a very personal look on the protagonists of our adventure.

Today I used my new 360° camera the first time. And I already learned a bit. To avoid shadows on the complete frame I will probably use it, when the sun stands the highest, or it isn’t there at all.

Overall it was a great day. I am running a bit behind with interviews, but I leave that for the days, when weather forbids us from going outside. The pictures today were fantastic and I am optimistic for the days to come.

Thank you very much for your help and attention.

Willie Schumann, Journalist, Crew 184

Copyright © The Mars Society. All rights reserved. | Main Site